OpenStudy (anonymous):

Given the equations E=1+3a,F=2a-b and G=ab-8.Find the possible values of a and of b so that E=2F=G @rational @phi

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

What is preventing you from parsing it and solving it? E=2F=G MEANS E = G -- Set that up. E = 2F -- Set that up. 2F = G -- Set that up. Once you get a look at all three, solve away.

OpenStudy (anonymous):

\[1+3a=ab-8\]\[1+3a=4a-2b\]\[4a-2b=ab-8\]

OpenStudy (anonymous):

a=-2,9 and b=-3/2,4

OpenStudy (anonymous):

Thnx @tkhunny